OpenStudy (anonymous):

What is the eccentricity of the ellipse if its major axis is twice its minor axis?

OpenStudy (anonymous):

eccentricity is M/m... so if M = 2m, then the eccentricity would be \[\frac{M}{m}=\frac{2m}{m}=?\]